# Merger History of Clustered Primordial Black Holes

### v2

@@ VII Conclusion
The paper investigated the influence of the initial clustering of PBHs on the dynamics of their mergers. It was shown that a significant part of the binaries are characterized by sufficiently wide orbits and small angular momenta, and therefore they are perturbed in the dense environment of the cluster. This process leads to an effective “elimination” of such binaries from the merger process due to a significant increase of their coalescence time or even destruction. At the same time, rare close binaries are more likely to avoid perturbations and therefore continue to merge. Such a change in the distribution of the PBH binaries orbital parameters causes a suppression of the merger rate starting from the time $t_{\rm sup}$ determined by Eq. (20), these results are shown in Fig. 2. In the epoch of active binaries perturbation, the merger rate scales with cosmological time as $\mathcal{R}\propto t^{-45/14}$ (in accordance with Eq. (V)). Which differs significantly from the case of a Poisson (initially non-clustered) distribution of PBHs in the space $\mathcal{R}\propto t^{-34/37}$. The study of the redshift evolution of black holes merger rate is one of the target for the next generation of gravitational wave detectors. Therefore, future observations will be able not only to verify the existence of PBHs, but also to reconstruct the features of their initial spatial distribution, which will make it possible to identify the most plausible scenarios for their formation. The absence of GW signals at high redshifts, in turn, will impose strict constraints on the parameters of PBHs formation models.
On the other hand, if direct detection of GW transients from black hole mergers at large redshifts turns out to be observationally difficult, their overall contribution may still be significant as the integral effect on the gravitational wave energy in the Universe. Therefore, in Section VI we analyze the impact of our results for the stochastic GW background. We demonstrate that a rapid growth in the black hole merger rate with redshift causes spectral features in the GW energy density, as illustrated in Fig. 4. Specifically, when the merger rate is scaled with time as $\mathcal{R}\propto t^{-45/14}$, the stochastic GW background depends on the frequency $\Omega_{\rm gw}\propto\nu^{-65/28}$ in the frequency range above $\nu_{\rm sup}$ given by Eq. (32). This is radically different from the traditional expectation of $\Omega_{\rm gw}\propto\nu^{2/3}$ for black hole mergers. The obtained results provide an additional possibility for experimentally testing PBH clustering scenarios.
@@ Acknowledgment
The work was funded by the Ministry of Science and Higher Education of the Russian Federation, Project ”New Phenomena in Particle Physics and the Early Universe” FSWU-2023-0073

### v3

@@ VII Conclusion
The paper investigated the influence of the initial clustering of PBHs on the dynamics of their mergers. It was shown that a significant part of the binaries are characterized by sufficiently wide orbits and small angular momenta, and therefore they are perturbed in the dense environment of the cluster. This process leads to an effective “elimination” of such binaries from the merger process due to a significant increase of their coalescence time or even destruction. At the same time, rare close binaries are more likely to avoid perturbations and therefore continue to merge. Such a change in the distribution of the PBH binaries orbital parameters causes a suppression of the merger rate starting from the time $t_{\rm sup}$ determined by Eq. (20), these results are shown in Fig. 2. In the epoch of active binaries perturbation, the merger rate scales with cosmological time as $\mathcal{R}\propto t^{-45/14}$ (in accordance with Eq. (V)). Which differs significantly from the case of a Poisson (initially non-clustered) distribution of PBHs in the space $\mathcal{R}\propto t^{-34/37}$. The study of the redshift evolution of black holes merger rate is one of the target for the next generation of GW detectors. Therefore, future observations will be able not only to verify the existence of PBHs, but also to reconstruct the features of their initial spatial distribution, which will make it possible to identify the most plausible scenarios for their formation. The absence of GW signals at high redshifts, in turn, will impose strict constraints on the parameters of PBHs formation models.
On the other hand, if direct detection of GW transients from black hole mergers at large redshifts turns out to be observationally difficult, their overall contribution may still be significant as the integral effect on the GW energy in the Universe. Therefore, in Section VI we analyze the impact of our results for the stochastic GW background. We demonstrate that a rapid growth in the black hole merger rate with redshift causes spectral features in the GW energy density, as illustrated in Fig. 4. Specifically, when the merger rate is scaled with time as $\mathcal{R}\propto t^{-45/14}$, the stochastic GW background depends on the frequency $\Omega_{\rm gw}\propto\nu^{-65/28}$ in the frequency range above $\nu_{\rm sup}$ given by Eq. (32). This is radically different from the traditional expectation of $\Omega_{\rm gw}\propto\nu^{2/3}$ for black hole mergers. We summarize our main findings relating to the scaling of the merger rate and the GW background in Table 1. The obtained results provide an additional possibility for experimentally testing PBH clustering scenarios.
@@ Acknowledgment
The work was funded by the Ministry of Science and Higher Education of the Russian Federation, Project “New Phenomena in Particle Physics and the Early Universe” FSWU-2023-0073. I am grateful to the anonymous reviewers for their constructive comments and recommendations, which helped improve the paper.
Appendix A: Mergers of perturbed binaries
To ensure completeness in Sec. V, this appendix outlines considerations relating to the merging of perturbed binaries. First of all, it is necessary to separate hard binaries from soft ones in the entire population of PBH binaries. For this purpose, it is sufficient to consider the distribution of distances between two PBHs $dP=3x^{2}/\overline{x}^{3}\,dx$ and integrate it to $x(a_{h})$:
$P_{h}\equiv P(a<a_{h})=\left(\frac{a_{h}\delta_{\rm PBH}}{\overline{x}}\right)^{3/4}$
$\simeq 2\times 10^{-3}\delta_{\rm PBH}\left(\frac{m}{M_{\odot}}\right)^{1/2}\left(\frac{\sigma}{10\,\rm km\,s^{-1}}\right)^{-3/2}.$ (33)
The relationship between $x$ and the major semiaxis $a$ is determined by Eq. (11), and $a_{h}$ is the major semiaxis separating hard binaries from soft ones (16).
Now, following the explanation in Ref. Heggie and Hut (2003), we will take a qualitative look at the dynamics of hard binaries in clusters. The change in the energy of a hard binary resulting from a single scattering event involving a PBH is comparable in magnitude to its binding $\Delta E\sim|\varepsilon|=Gm^{2}/2a$. The conservation of energy during three-body scattering in the rest frame of the barycenter system reads
$\frac{\mu v^{2}}{2}+\varepsilon=\frac{\mu v^{\prime 2}}{2}+\varepsilon(1+\Delta),$ (34)
here, $v$ and $v^{\prime}$ are the relative velocities at infinity of the binary and single PBH, respectively, before and after scattering, $\mu=m/3$ is the reduced mass and $\Delta=(\varepsilon^{\prime}/\varepsilon-1)\simeq 1$ is the relative change in binding energy of binary after scattering. In the case of a sufficiently hard binary, i.e. when $|\varepsilon|\gg\mu v^{2}$, the relative velocity after scattering is given by $v^{\prime}\simeq\sqrt{-3\,\varepsilon\Delta/m}$ and the velocity of the binary will be $v^{\prime}_{b}=v^{\prime}/3$. To obtain the change in velocity relative to the whole cluster, it is necessary to add the velocity of the triple system, but this is negligible within the limits of a sufficiently hard binary. Thus, the fate of the binary system is as follows: due to three-body scattering, the binary becomes increasingly bound until it ultimately acquires sufficient recoil to leave the cluster. Also, since the recoil velocity depends solely on the binding energy, we assume as a first approximation that the perturbed binaries have approximately the same binding energy $\varepsilon=km\sigma^{2}/2$ and merge outside the cluster. Moreover, due to perturbations, their initial angular momentum distribution is strongly mixed (and likely approaches to thermal one), so we assume the distribution of binary parameters to be as follows:
$dP_{\rm per}=P_{h}(n+1)\delta(a-a_{*})j^{n}\,dadj,$ (35)
where $a_{*}=Gm/k\sigma^{2}$, $\delta(a-a_{*})$ is the delta function and $k\gg 1$ (for globular star clusters $k\simeq 100$ typically Heggie and Hut (2003); Binney and Tremaine (2008)). Note that the case $n=1$ corresponds to thermal distribution. The fraction of binaries merging by time $t$ is defined as $P_{\rm per}=\int_{0}^{j_{*}}dP_{\rm per}=P_{h}j_{*}^{n+1}$, where
$j_{*}=\left(\frac{170\,G^{3}m^{3}t}{3\,c^{5}a_{*}^{4}}\right)^{1/7}$ (36)
is obtained using the merger time (13). Here, for simplicity, we consider the case that perturbed binaries formed in a enough young Universe. In other words, the time difference between the formation epoch and the merger of these binaries corresponds to cosmological time. However, it is clear that this assumption may be inaccurate, particularly for binaries that initially formed a very close system (their characteristic perturbation time is large). The merger rate is determined by a similar expression to that given in Eq. (25)
$\mathcal{R}_{\rm per}=\frac{f\Omega_{\rm DM}}{m}\frac{3H_{0}^{2}}{8\pi G}\frac{dP_{\rm per}}{dt}$
$=\frac{\rho_{\rm DM}f\delta_{\rm PBH}}{m}\frac{n+1}{7t}\left(\frac{Gm^{2/3}\rho_{\rm eq}}{2\sigma^{2}}\right)^{3/4}\left(\frac{170k^{4}\sigma^{8}t}{3c^{5}Gm}\right)^{(n+1)/7}.$ (37)
For this result to be consistent, a transition to the unperturbed binary merger regime, defined by Eq. (V), must occur, but this analysis is beyond the scope of this paper. By numerically estimating the merger rate (Appendix A: Mergers of perturbed binaries) for some parameters, we obtain
$\mathcal{R}_{\rm per}$ $\simeq\dfrac{3.4\times 10^{4}}{\rm Gpc^{3}yr}f\delta_{\rm PBH}\left(\dfrac{k}{100}\right)^{8/7}\left(\dfrac{m}{10\,M_{\odot}}\right)^{-11/14}\left(\dfrac{\sigma}{10\,\rm km\,s^{-1}}\right)^{11/14}\left(\dfrac{t}{t_{0}}\right)^{-5/7}\quad{\rm{for}}\,n=1,$
$\mathcal{R}_{\rm per}$ $\simeq\dfrac{6.6\times 10^{4}}{\rm Gpc^{3}yr}f\delta_{\rm PBH}\left(\dfrac{k}{100}\right)^{4/7}\left(\dfrac{m}{10\,M_{\odot}}\right)^{-9/14}\left(\dfrac{\sigma}{10\,\rm km\,s^{-1}}\right)^{-5/14}\left(\dfrac{t}{t_{0}}\right)^{-6/7}\quad{\rm{for}}\,n=0,$ (38)
where we consider two illustrative examples for angular momentum distribution. However, we note that depending on the velocity dispersion $\sigma$, we note that the the merger rate can vary by several orders of magnitude as a function of $n$, as shown in Fig. 5. We also note that, for all reasonable values of the parameter $n$ in Eq. (35), the merger rate of perturbed binaries is a slow function of cosmological time. Therefore, their merging does not lead to the appearance of spectral features in the stochastic GW background, similar to those discussed in Sec. VI.
Figure 5: The current merger rate as a function of the angular momentum distribution parameter $n$ in Eq. (35). Different curves correspond to different velocity dispersions $\sigma$.
Appendix B: Mergers of stellar origin black holes
Figure 6: The normalized merger rate of binary black holes of different origins. The red dashed line corresponds to astrophysical black holes defined by Eq. (Appendix B: Mergers of stellar origin black holes). The blue solid and dash-dotted lines show the mergers of unperturbed PBHs in the case of their initial clustering, which are scaled as $\mathcal{R}\propto t^{-45/14}$ and $\mathcal{R}\propto t^{-61/28}$, respectively (see Sec. V for details). The dotted purple line represents the merging of perturbed PBH binaries with the thermal distribution of angular momentum $\mathcal{R}\propto t^{-5/7}$ ($n=1$ in Eqs. (35) and (Appendix A: Mergers of perturbed binaries)). The black solid line shows mergers of PBHs without any initial clustering in their spatial distribution $\mathcal{R}\propto t^{-34/37}$.
This appendix considers a more traditional source of GW events: mergers of astrophysical black holes. Here, we will follow the analysis of Raidal et al. (2019); Mukherjee and Silk (2021). The astrophysical black hole merger rate should follow the rate of star formation, which is fitted by the empirical formula Madau and Dickinson (2014)
$\mathcal{R}_{\rm{SFR}}(z)\propto\frac{(1+z)^{2.7}}{1+\big((1+z)/2.9\big)^{5.6}}.$ (39)
Since the formation of stellar binaries, the subsequent formation of black holes, and their mergers are separated in time, a time delay must be introduced between these processes. Similar to Ref. Belczynski et al. (2016), we assume that the delay time distribution is $P_{\rm del}\propto t^{-1}$ and zero otherwise. Then the dependence of the merger rate of astrophysical black holes on redshift is determined by
$\mathcal{R}_{\rm{ABH}}(z)\propto\int_{z}^{\infty}dz_{f}\,\frac{dt_{f}}{dz_{f}}\,dt_{\rm del}\,\mathcal{R}_{\rm{SFR}}(z_{f})$
$\times P_{\rm del}(t_{\rm del})\delta\Big(t(z)-t(z_{f})-t_{\rm del}\Big),$ (40)
where $z_{f}$ is the redshift of the formation of the stellar binary and $dt/dz=(H(z)(1+z))^{-1}$.
Figure 6 shows the temporal evolution of the merger rate of astrophysical and primordial black holes, taking various clustering scenarios into account. As can be seen, mergers of PBHs inevitably dominate astrophysical black holes at large redshifts. It is widely known that this idea opens up the possibility of testing the existence of PBHs by observing GW transients at large cosmological distances, even when $f_{\rm PBH}\ll 1$. However, it should be noted that observing black hole mergers at high redshifts will be challenging for GW astronomy in the near future.